Solve the system of equations by the substitution method.
-6y=x+21
3x+5y=11
Please Help.

Answer by stanbon(75887)

Solve the system of equations by the substitution method.
-6y=x+21
3x+5y=11
----
Solve the top equation for "x"::
x = -6y-21
----
Substitute that into the bottom equation and solve for "y"::
3(-6y-21) + 5y = 11
-18y - 63 + 5y = 11
-13y - 63 = 11
13y = -63-11
y = -74/13 = -5.69
----------
Solve for "x":
x = -6y-21
---
x = -6(-74/13) - 21
x = 444/13 - 21 = 13.15
---------------
Cheers,
Stan H.
-------------